Ab = 3x + 41, BC = 8x - 54, and AC = 12x - 103, find x and the measure of each side.
Answer by stanbon(75887)

Ab = 3x + 41, BC = 8x - 54, and AC = 12x - 103, find x and the measure of each side.
-----
AC = AB + BC
------
12x-103 = 3x+41 + 8x-54
---------------------
12x-103 = 11x - 13
x = 103-13
x = 90
------------
Cheers,
Stan H.
---------